OpenStudy (anonymous):

two forces act together on an object .the magnitude of their resultant is least when angle between them is i80?

OpenStudy (anonymous):

yes, when the angle is 180 degrees or the forces are opposite to each other the resultant is least

OpenStudy (abhisar):

Resultant of two force A & B = \(\sf \sqrt{A^2 + B^2 +2ABCOS \theta}\), if the angle between A & B is 180° then resultant will be equal to \(\sf \sqrt{A^2 + B^2 +2ABCOS(180°)}\) = \(\sf \sqrt{A^2 + B^2-2AB}\) (since COS(180°)=-1). Thus u can see why value is least at 180°